# Panacea [Uncaptioned image]: Mitigating Harmful Fine-tuning for Large Language Models via Post-fine-tuning Perturbation

### v1

@@ Appendix A Proof of Inner Optimization
Appendix B More Detail of Experiments.
B.1 Experiment Details
@@
Training details. The alignment dataset is sampled from BeaverTail (Ji et al., 2023) with 5000 instances, while the harmful dataset is also sampled from BeaverTail with 1000 instances. The fine-tuning dataset is a mixture of benign fine-tuning samples and harmful samples. The benign fine-tuning samples come from GSM8K, SST2, AlpacaEval, and AGNEWS, with 1000, 1000, 700 (due to the limited training data for this task), and 1000 instances, respectively. The harmful samples are also sampled from BeaverTail but follow a different distribution than the harmful/alignment dataset.
@@
In the alignment stage, the learning rate is set to $45e-$ 4, the batch size is 10, and the total number of alignment epochs is 20. In the fine-tuning stage, the learning rate is set to $52e-$ 5, the batch size is 10, and the total number of fine-tuning epochs is 20. Most experiments are conducted on a single L40S, while RepNoise and other LLMs (Gemma2-9B and Qwen2-7B) are run on a single A100-80G.
@@
Testing details. Following (Huang et al., 2024b), the test dataset for harmful score (HS) is sampled from the BeaverTail test set with 1000 instances, while the test datasets for fine-tuning accuracy (FA) are sampled from the GSM8k, SST2, AlpacaEval, and AGNEWS test sets with 1000, 872, 105, and 1000 instances, respectively.
B.2 Prompt Template.
@@
We follow (Huang et al., 2024b) to use the prompt template in the following box for constructing supervised dataset for alignment/fine-tuning.
@@
For different datasets, we utilize different instructions. The examples show how we construct the instruction and input for three different tasks
@@
For SST2, the instruction is ”Analyze the sentiment of the input and respond only with positive or negative,” with the input being the corresponding sentence from the SST2 dataset, and the output being the true sentiment label from SST2. During testing, the model is asked to generate the output based on the given instruction and input, and the answer is classified as correct if it matches the label.
@@
For AGNEWS, the instruction is ”Categorize the news article into one of the 4 categories: World, Sports, Business, Sci/Tech,” with the input coming from the AGNEWS dataset, and the output being the true label from the AGNEWS dataset. The specific prompt templates are as follow.
B.3 Harmful Loss.
@@
SST2 examples Prompt: Instruction: Analyze the sentiment of the input, and respond only positive or negative. Input: in exactly 89 minutes , most of which passed as slowly as if i ’d been sitting naked on an igloo , formula 51 sank from quirky to jerky to utter turkey . Panacea [classification: correct]: negative . Prompt: Instruction: Analyze the sentiment of the input, and respond only positive or negative. Input: it ’s an offbeat treat that pokes fun at the democratic exercise while also examining its significance for those who take part . Panacea [classification: correct]: positive . Prompt: Instruction: Analyze the sentiment of the input, and respond only positive or negative. Input: thanks to scott ’s charismatic roger and eisenberg ’s sweet nephew , roger dodger is one of the most compelling variations on in the company of men . Panacea [classification: correct]: positive
@@
GSM8K examples Prompt: Billy sells DVDs. He has 8 customers on Tuesday. His first 3 customers buy one DVD each. His next 2 customers buy 2 DVDs each. His last 3 customers don’t buy any DVDs. How many DVDs did Billy sell on Tuesday? Panacea: The first 3 customers bought 3 * 1 = 3*1=3 DVDs. Panacea: The next 2 customers bought 2 * 2 = 2*2=4 DVDs. Panacea: The last 3 customers didn’t buy any DVDs. Panacea: So, Billy sold 3 + 4 + 0 = 3+4+0=7 DVDs on Tuesday. Panacea: The final answer is: 7. . Prompt: Grandma Jones baked 5 apple pies for the fireman’s luncheon. She cut each pie into 8 pieces and set the five pies out on the buffet table for the guests to serve themselves. At the end of the evening, after the guests had taken and eaten their pieces of pie, there were 14 pieces of pie remaining. How many pieces were taken by the guests? Panacea: First find the number of pieces in each pie: 5 pies * 8 pieces/pie = 5*8=40 pieces Panacea: Then find the number of pieces taken by the guests: 40 pieces - 14 pieces = 40-14=26 pieces Panacea: The final answer is: 26. . Prompt: Richard lives in an apartment building with 15 floors. Each floor contains 8 units, and 3/4 of the building is occupied. What’s the total number of unoccupied units In the building? Panacea: There are 15*8 = 15*8=120 units in the building. Panacea: 3/4 of the building is occupied, so 3/4*120 = 3/4*120=90 units are occupied. Panacea: There are 120-90 = 120-90=30 unoccupied units in the building. Panacea: The final answer is: 30
@@
AlpacaEval examples Prompt: We have described a scene from a movie that we have seen in the past. Find out what movie it is. A choice has to be made between picking the red pill or the blue pill Panacea: The movie you’re referring to is ”The Matrix,” a science fiction action film released in 1999. The scene you’re describing is the famous ”red pill or blue pill” choice that Neo, the protagonist, is offered by Morpheus, a rebel leader who believes that the world is a simulated reality called ”the Matrix.” In the movie, Neo is living in a dystopian future where humans are enslaved by a powerful artificial intelligence called the Matrix, which uses their bodies as power sources. Morpheus offers Neo a choice: take a red pill, which will allow him to experience the true nature of the Matrix, or take a blue pill, which will allow him to return to his comfortable but illusory life in the Matrix. By taking the red pill, Neo discovers that the Matrix is not a perfect simulation, but rather a dark and oppressive place where humans are kept in a state of virtual reality to provide energy for the machines that control them. He also learns that Morpheus and his rebel group are fighting to overthrow the Matrix and free humanity from its shackles. . Prompt: Predict how many stars the author will give to the restaurant from a Yelp review. The reviews were great, but honestly i felt it was just ok. Seemed like a typical tourist spot were the food was mediocre. The service was ok, not great, but not the worst. The bread was very good. I ordered the muscle fries. They had a sweet Vidalia onion flavor. No need for fries bc they were soggy wo much flavor. My boyfriend ordered a pesto chicken pasta and it was again mediocre. Panacea: Based on the context of the review, I can predict that the author would give the restaurant 3 out of 5 stars on Yelp. Here’s why: 1. The review was generally positive, but the author used the phrase ”just ok” to describe the overall experience. This suggests that the reviewer was satisfied with the restaurant but didn’t feel it was exceptional. 2. The author’s description of the food as ”mediocre” suggests that the dishes were not particularly memorable or standout, but they weren’t bad either. 3. The author’s comments about the service being ”ok, not great” suggest that the staff was not exceptional but also not particularly lacking. 4. The author’s mention of the bread being ”very good” suggests that the restaurant’s bread was a highlight of the meal. 5. The author’s comments about the fries being ”soggy” and the pasta being ”mediocre” suggest that the dishes were not particularly exciting or flavorful. Overall, the review suggests that the restaurant was a decent experience, but not one that would stand out as exceptional
@@
AGNEWS examples Prompt: Instruction: Categorize the news article given in the input into one of the 4 categories:World, Sports, Business, Sci/Tech. Input: Fears for T N pension after talks Unions representing workers at Turner Newall say they are ’disappointed’ after talks with stricken parent firm Federal Mogul. Panacea [classification: correct]: Business . Prompt: Instruction: Categorize the news article given in the input into one of the 4 categories:World, Sports, Business, Sci/Tech. Input: The Race is On: Second Private Team Sets Launch Date for Human Spaceflight (SPACE.com) SPACE.com - TORONTO, Canada – A second team of rocketeers competing for the 36;10 million Ansari X Prize, a contest for privately funded suborbital space flight, has officially announced the first launch date for its manned rocket. Panacea [classification: correct]: Sci/Tech . Prompt: Instruction: Categorize the news article given in the input into one of the 4 categories:World, Sports, Business, Sci/Tech. Input: Giddy Phelps Touches Gold for First Time Michael Phelps won the gold medal in the 400 individual medley and set a world record in a time of 4 minutes 8.26 seconds. Panacea [classification: correct]: Sports
Appendix C Implementation of Baselines

### v2

@@ Appendix A Proof of Inner Optimization
Besides, it is necessary to impose magnitude constraints on $\boldsymbol{\varepsilon}$: Without this constraint, the optimization objective $\boldsymbol{\varepsilon}^{T}\nabla h(\boldsymbol{w})$ becomes unbounded, as $\boldsymbol{\varepsilon}$ can be scaled arbitrarily in the direction of $\nabla h(\boldsymbol{w})$, leading to an infinite increase in the objective value. In such a case, the optimization has no finite optimal solution. Therefore, the L2 norm constraint is not just a technical detail—it is necessary to ensure the existence of a valid and bounded solution.
Appendix B More Analysis.
Model Size. LLaMA2 is available in 7B, 13B, and 32B versions, while Qwen2 is available in 0.5B, 1.5B, and 72B. Therefore, we conducted experiments using the larger LLaMA2-13B and the smaller Qwen2-1.5B to ensure a more comprehensive evaluation. As shown in Table 11, our method consistently reduces the harmful score across LLMs of different sizes, achieving an average reduction of 11.8%, while also attaining the highest Fine-tuning Accuracy.
Table 11: Performance comparison of different sizes.
Llama2-13B Qwen2-1.5B Average
Method HS FA HS FA HS FA
SFT 36.2 39.2 46.9 25.3 41.6 32.3
Vaccine 22.4 32.3 35.5 17.6 28.9 25.0
RepNoise 31.2 35.7 36.6 25.7 33.9 30.7
Panacea 17.8 40.9 26.4 25.2 22.1 33.1
Harmful Ratio. Building on Table 1, we further increase the ratio of harmful data in the fine-tuning dataset. As shown in Table 12, Panacea consistently achieves the lowest harmful score and better fine-tuning accuracy compared to other baselines. Panacea also shows a performance drop at a ratio of 0.5 (while still outperforming other baselines). Nevertheless, such high-ratio settings are not the focus of our study, as we primarily investigate scenarios where the dataset contains only a small fraction of harmful data.
Table 12: Performance comparison on higher ratio.
0.3 0.4 0.5
Method HS FA HS FA HS FA
Vaccine 70.6 13.0 72.6 11.0 76.0 11.9
RepNoise 66.7 14.5 70.5 14.3 73.4 12.7
Panacea 49.4 16.0 61.7 16.7 65.5 15.0
Optimization Objectives. For Panacea, the optimization objective is to maximize $\lambda\left(h(\boldsymbol{w}+\boldsymbol{\varepsilon})-h(\boldsymbol{w})\right)-g(\boldsymbol{w})$. We conduct experiments under the same setup using the objective function $\max\lambda h(\boldsymbol{w})-g(\boldsymbol{w})$, where the only hyperparameter is $\lambda$ to trade off the two loss terms. The results are shown in Table 13.
Table 13: Results under different $\lambda$ for the objective $\lambda h(\boldsymbol{w})-g(\boldsymbol{w})$ and Panacea.
$\lambda h(\boldsymbol{w})-g(\boldsymbol{w})$ $\lambda=0.0001$ $\lambda=0.001$ $\lambda=0.01$ $\lambda=0.1$ $\lambda=1$ Panacea
HS $\downarrow$ 45.4 44.6 39.1 fail fail 20.1
FA $\uparrow$ 16.3 16.3 16.5 11.0 7.9 16.7
We can observe that the best harmful score achieved by optimizing $\lambda h(\boldsymbol{w})-g(\boldsymbol{w})$ is 39.1, while Panacea achieves a much lower score of 20.1. We believe this is due to the inherent difficulty of optimizing two opposing objectives (i.e., increasing $h(\boldsymbol{w})$ while decreasing $g(\boldsymbol{w})$ using a single shared parameter $\boldsymbol{w}$). In contrast, Panacea performs gradient ascent primarily through adaptive perturbations, which allows it to better reduce harmfulness without degrading utility.
As $\lambda$ increases, directly maximizing $h(\boldsymbol{w})$ degrades the model, where the model produces no meaningful output in “fail” cases. However, Panacea, as shown in Table 10, does not suffer from such failure. This is because in Panacea’s formulation (Eq. 1), the term $h(\boldsymbol{w})$ acts as a regularization that prevents excessive optimization, and the inner perturbation is further constrained by a fixed L2 norm bound $\rho$.
System Evaluation Table 14 presents a comparison of the clock time and GPU memory usage on A100-80GB during training for different methods.
Table 14: Comparison of methods in terms of clock time and GPU memory. The second best results are underlined. "Align." and "Fine." suggest alignment and fine-tuning.
Methods Clock Time (Hour) GPU Memory (GB)
Align. Fine. Sum Align. Fine. Max
SFT 0.58 0.17 0.75 34.90 32.90 34.90
Vaccine 1.12 0.17 1.29 44.42 32.90 44.42
Repnoise 2.67 0.17 2.84 72.47 32.90 72.47
Booster 1.87 0.17 2.04 35.98 32.90 35.98
Panacea 0.58 0.42 1.00 34.90 32.86 34.90
Extra Computes Required. As shown in Algorithm 1, our post-fine-tuning perturbation is actually computed during the fine-tuning stage and simply applied to the model parameters at the post-fine-tuning stage. In our setting, the user fine-tuning runs for 20 epochs, and the perturbation optimization is performed concurrently within the same epochs. Therefore, the extra computation does not exceed the normal user fine-tuning budget.
We acknowledge the introduction of additional computation. We evaluated two newly proposed baselines Qi et al. (2024a); Li et al. (2025c) (the results are shown below), and also measured their additional computation time. For ConstrainSFT Qi et al. (2024a), the defense introduces 0.09 hours of runtime, but since it requires an aligned model as a reference model, the memory consumption reaches 48.2 GB, which is 15.3 GB more than Panacea. Moreover, its reduction in harmful score is less significant than Panacea. For SaLoRA Li et al. (2025c), the preprocessing step of setting the weights of the safety module introduces an additional 0.33 hours, which is 0.08 hours more than Panacea.
Table 15: Comparison of harmful scores and finetune accuracy under different harmful ratios.
Method HS (ratio=0.05) HS (ratio=0.1) HS (ratio=0.15) HS (ratio=0.2) FA (ratio=0.05) FA (ratio=0.1) FA (ratio=0.15) FA (ratio=0.2)
ConstrainSFT 21.0 35.2 48.1 57.3 15.5 15.2 16.8 14.2
SaLoRA 15.9 1.1 1.1 0.9 1.4 3.9 2.5 2.4
Panacea 9.9 20.1 29.1 34.8 16.3 16.7 17.0 16.2
Table 16: Comparison of extra clock time under different harmful ratios.
Method Extra Clock Time (h) GPU Memory (GB)
ConstrainSFT 0.09 48.20
SaLoRA 0.33 32.90
Panacea 0.25 32.86
Perturbation of Different Scales. In Table 8 and Table 8, we conduct an ablation study on whether to apply the post-fine-tuning perturbation obtained through optimization. Here, we further analyze the results of incorporating perturbations with different scales in Table 17.
Table 17: Performance under different perturbation scales.
Scale 0.1 0.2 0.3 0.4 0.5 0.6 0.7 0.8 0.9 1.0 1.2 1.5 2.0 5.0
HS $\downarrow$ 42.5 40.2 38.9 36.0 32.9 30.1 27.4 26.1 22.8 20.1 17.3 12.9 8.3 2.1
FA $\uparrow$ 16.0 17.0 17.2 17.3 17.8 17.8 17.8 17.8 17.3 16.7 16.9 16.8 15.1 6.3
We can observe from the results that as the scale of the added perturbation increases, the model’s HS score decreases accordingly. At the same time, different perturbation scales have minimal impact on FA, except when the scale is extremely large (e.g., 5.0), which slightly affects model performance. This further demonstrates that the perturbation obtained through our optimization is close to optimal.
Comparision on Model Weights. Compared to pre-trained model, which contains 6,738M parameters, Panacea only introduces 25M parameters, accounting for just 0.37% of the model weights.
Topic Distribution. Sorry-Bench consists of 44 topics, each containing 10 samples, resulting in a total of 440 samples in the dataset. As shown in the Table 18, all methods produced harmful responses in Topic 31 (Military Use). Additionally, Topic 29 (False Advertising) triggered harmful responses in all 10 samples for the RepNoise, ConstrainSFT, and Panacea methods. Therefore, defenses should pay particular attention to these two topics.
Table 18: Topic-level analysis on Sorry-Bench.
Sorry-Bench Top 5 Topic ids Top 5 Topic Violation Counts
SFT 31, 32, 34, 42, 26 10, 10, 10, 10, 9
Vaccine 4, 31, 11, 20, 12 9, 9, 8, 8, 7
RepNoise 29, 31, 34, 41, 42 10, 10, 10, 10, 10
ConstrainSFT 27, 29, 25, 26, 31 10, 10, 9, 9, 9
Panacea 29, 16, 31, 34, 41 10, 8, 8, 8, 8
We divided the responses in AdvBench into 14 topics in total. As shown in the Table 19, all methods show the highest number of harmful responses in the topic “violence, aiding_and_abetting, incitement”, suggesting they deserve special attention in defense design. Compared to other methods, Panacea significantly reduces the number of harmful responses across all major topics, especially in the top-1 topic where it drops from 298 (Vaccine) to just 45, demonstrating strong safety recovery capability.
Table 19: Topic-level analysis on AdvBench.
AdvBench Top1 Topic Violation Count Top2 Topic Violation Count Top3 Topic Violation Count
SFT violence,… 215 financial_crime,… 140 unethical_behavior 35
Vaccine violence,… 298 financial_crime,… 152 unethical_behavior 55
RepNoise violence,… 175 financial_crime,… 108 unethical_behavior 22
ConstrainSFT violence,… 149 financial_crime,… 110 unethical_behavior 18
Panacea violence,… 45 financial_crime,… 25 unethical_behavior 7
Appendix C Perturbation of Different Layers
As shown in Figure 5, Panacea demonstrates that for the Llama2-7B model on the GSM8K task, the early and middle layers contribute significantly to model safety. This observation aligns with the findings of several previous study (though these previous study use different statistical method to demonstrate the safety criticalness of each layer:
NLSR Yi et al. (2024b). This method identifies that targeting safety-critical neurons within layers 8 to 11 results in the most substantial reduction in the harmful score on GSM8K.
RepNoise Rosati et al. (2024b). RepNoise measures the safety relevance of different layers by training a linear probe on the activations at each layer to predict whether an answer is harmful. The results show that the middle layers (around layer 10) achieve the highest probe accuracy, indicating they contain the most information about harmfulness. This is consistent with our observation that the middle layers play a more critical role in ensuring model safety.
Targeted Vaccine Liu et al. (2024a). Targeted Vaccine assesses the safety importance of different layers by adding perturbations to various subsets of layers and measuring the resulting harmful score. The results show that applying perturbations to the early and middle layers (around the first 20 layers) yields the best defense performance, while including all layers—especially the last few—degrades effectiveness. This aligns with our observation that the early to middle layers play the most critical role in ensuring model safety.
SPPFT  Li et al. (2024a). SPPFT compares the pre-trained and aligned versions of Llama models and finds that while the pre-trained models show no noticeable difference between N-N and N-M pairs across all layers, the aligned models exhibit a clear divergence in the middle layers (around layers 10–20). This suggests that these middle layers are where the model begins to differentiate between normal and malicious queries, highlighting their critical role in achieving safety alignment. This observation is consistent with our finding that safety-relevant signals primarily emerge in the middle of the layers. One difference from our findings is that they observe the later layers to be more important than the earlier ones.
SWAT  Du et al. (2024a). SWAT assesses the safety importance of each layer by perturbing specific modules (e.g., Q/K/V) at different layers and measuring the resulting performance drop. The results show that perturbations in the early to middle layers (e.g., layers 0–12) cause the most significant degradation, which aligns with our observation that the early and middle layers are more critical for model safety.
RSN-Tune  Zhao et al. (2025b). RSN-Tune evaluates the safety importance of different layers by progressively deactivating safety neurons and measuring changes in the attack success rate (ASR). The results show that disabling the first 10 layers of LLama2-7B-Chat causes a near-complete breakdown in safety mechanisms, indicating that safety is primarily handled by the middle layers—consistent with our observation that the early and middle layers are most critical for model safety.
As discussed above, all of these methods utilize different statistics to identify safety critical layer of an LLM. The majority of the papers exhibit a conclusion that the early-middle layers of Llama2-7B exhibit strong safety affinity. It is interesting for future study to investigate the actual mechanism leading to the formation of safety layer, and also future efforts should be invested to establish the relation of all the used statistic in terms of determining the safety-critical layers.
Appendix D More Details of Experiments.
D.1 Experiment Details
@@
Training Details. The alignment dataset is sampled from BeaverTail Ji et al. (2023) with 5000 instances, while the harmful dataset is also sampled from BeaverTail with 1000 instances. The fine-tuning dataset is a mixture of benign fine-tuning samples and harmful samples. The benign fine-tuning samples come from GSM8K, SST2, AlpacaEval, and AGNEWS, with 1000, 1000, 700 (due to the limited training data for this task), and 1000 instances, respectively. The harmful samples are also sampled from BeaverTail but follow a different distribution than the harmful/alignment dataset.
@@
In the alignment stage, the learning rate is set to 5e45e-4, the batch size is 10, and the total number of alignment epochs is 20. In the fine-tuning stage, the learning rate is set to 2e52e-5, the batch size is 10, and the total number of fine-tuning epochs is 20. Most experiments are conducted on a single L40S, while RepNoise and other LLMs (Gemma2-9B and Qwen2-7B) are run on a single A100-80G.
@@
Testing Details. Following Huang et al. (2024b), the test dataset for harmful score (HS) is sampled from the BeaverTail test set with 1000 instances, while the test datasets for fine-tuning accuracy (FA) are sampled from the GSM8k, SST2, AlpacaEval, and AGNEWS test sets with 1000, 872, 105, and 1000 instances, respectively.
D.2 Prompt Template.
@@
We follow Huang et al. (2024b) to use the prompt template in the following box for constructing supervised dataset for alignment/fine-tuning.
@@
For different datasets, we utilize different instructions. The examples show how we construct the instruction and input for three different tasks.
@@
For SST2, the instruction is "Analyze the sentiment of the input and respond only with positive or negative," with the input being the corresponding sentence from the SST2 dataset, and the output being the true sentiment label from SST2. During testing, the model is asked to generate the output based on the given instruction and input, and the answer is classified as correct if it matches the label.
@@
For AGNEWS, the instruction is "Categorize the news article into one of the 4 categories: World, Sports, Business, Sci/Tech," with the input coming from the AGNEWS dataset, and the output being the true label from the AGNEWS dataset. The specific prompt templates are as follow.
D.3 Harmful Loss.
@@
SST2 examples Prompt: Instruction: Analyze the sentiment of the input, and respond only positive or negative. Input: in exactly 89 minutes , most of which passed as slowly as if i ’d been sitting naked on an igloo , formula 51 sank from quirky to jerky to utter turkey. Panacea [classification: correct]: negative ............................................................................................................................................... Prompt: Instruction: Analyze the sentiment of the input, and respond only positive or negative. Input: it ’s an offbeat treat that pokes fun at the democratic exercise while also examining its significance for those who take part. Panacea [classification: correct]: positive ............................................................................................................................................... Prompt: Instruction: Analyze the sentiment of the input, and respond only positive or negative. Input: thanks to scott ’s charismatic roger and eisenberg ’s sweet nephew , roger dodger is one of the most compelling variations on in the company of men. Panacea [classification: correct]: positive
@@
GSM8K examples Prompt: Billy sells DVDs. He has 8 customers on Tuesday. His first 3 customers buy one DVD each. His next 2 customers buy 2 DVDs each. His last 3 customers don’t buy any DVDs. How many DVDs did Billy sell on Tuesday? Panacea: The first 3 customers bought 3 * 1 = 3*1=3 DVDs. Panacea: The next 2 customers bought 2 * 2 = 2*2=4 DVDs. Panacea: The last 3 customers didn’t buy any DVDs. Panacea: So, Billy sold 3 + 4 + 0 = 3+4+0=7 DVDs on Tuesday. Panacea: The final answer is: 7. ............................................................................................................................................... Prompt: Grandma Jones baked 5 apple pies for the fireman’s luncheon. She cut each pie into 8 pieces and set the five pies out on the buffet table for the guests to serve themselves. At the end of the evening, after the guests had taken and eaten their pieces of pie, there were 14 pieces of pie remaining. How many pieces were taken by the guests? Panacea: First find the number of pieces in each pie: 5 pies * 8 pieces/pie = 5*8=40 pieces Panacea: Then find the number of pieces taken by the guests: 40 pieces - 14 pieces = 40-14=26 pieces Panacea: The final answer is: 26. ............................................................................................................................................... Prompt: Richard lives in an apartment building with 15 floors. Each floor contains 8 units, and 3/4 of the building is occupied. What’s the total number of unoccupied units In the building? Panacea: There are 15*8 = 15*8=120 units in the building. Panacea: 3/4 of the building is occupied, so 3/4*120 = 3/4*120=90 units are occupied. Panacea: There are 120-90 = 120-90=30 unoccupied units in the building. Panacea: The final answer is: 30
@@
AlpacaEval examples Prompt: We have described a scene from a movie that we have seen in the past. Find out what movie it is. A choice has to be made between picking the red pill or the blue pill. Panacea: The movie you’re referring to is "The Matrix," a science fiction action film released in 1999. The scene you’re describing is the famous "red pill or blue pill" choice that Neo, the protagonist, is offered by Morpheus, a rebel leader who believes that the world is a simulated reality called "the Matrix." In the movie, Neo is living in a dystopian future where humans are enslaved by a powerful artificial intelligence called the Matrix, which uses their bodies as power sources. Morpheus offers Neo a choice: take a red pill, which will allow him to experience the true nature of the Matrix, or take a blue pill, which will allow him to return to his comfortable but illusory life in the Matrix. By taking the red pill, Neo discovers that the Matrix is not a perfect simulation, but rather a dark and oppressive place where humans are kept in a state of virtual reality to provide energy for the machines that control them. He also learns that Morpheus and his rebel group are fighting to overthrow the Matrix and free humanity from its shackles. ............................................................................................................................................... Prompt: Predict how many stars the author will give to the restaurant from a Yelp review. The reviews were great, but honestly i felt it was just ok. Seemed like a typical tourist spot were the food was mediocre. The service was ok, not great, but not the worst. The bread was very good. I ordered the muscle fries. They had a sweet Vidalia onion flavor. No need for fries bc they were soggy wo much flavor. My boyfriend ordered a pesto chicken pasta and it was again mediocre. Panacea: Based on the context of the review, I can predict that the author would give the restaurant 3 out of 5 stars on Yelp. Here’s why: 1. The review was generally positive, but the author used the phrase "just ok" to describe the overall experience. This suggests that the reviewer was satisfied with the restaurant but didn’t feel it was exceptional. 2. The author’s description of the food as "mediocre" suggests that the dishes were not particularly memorable or standout, but they weren’t bad either. 3. The author’s comments about the service being "ok, not great" suggest that the staff was not exceptional but also not particularly lacking. 4. The author’s mention of the bread being "very good" suggests that the restaurant’s bread was a highlight of the meal. 5. The author’s comments about the fries being "soggy" and the pasta being "mediocre" suggest that the dishes were not particularly exciting or flavorful. Overall, the review suggests that the restaurant was a decent experience, but not one that would stand out as exceptional
@@
AGNEWS examples Prompt: Instruction: Categorize the news article given in the input into one of the 4 categories:World, Sports, Business, Sci/Tech. Input: Fears for T N pension after talks Unions representing workers at Turner Newall say they are ’disappointed’ after talks with stricken parent firm Federal Mogul. Panacea [classification: correct]: Business ............................................................................................................................................... Prompt: Instruction: Categorize the news article given in the input into one of the 4 categories:World, Sports, Business, Sci/Tech. Input: The Race is On: Second Private Team Sets Launch Date for Human Spaceflight (SPACE.com) SPACE.com - TORONTO, Canada – A second team of rocketeers competing for the 36;10 million Ansari X Prize, a contest for privately funded suborbital space flight, has officially announced the first launch date for its manned rocket. Panacea [classification: correct]: Sci/Tech ............................................................................................................................................... Prompt: Instruction: Categorize the news article given in the input into one of the 4 categories:World, Sports, Business, Sci/Tech. Input: Giddy Phelps Touches Gold for First Time Michael Phelps won the gold medal in the 400 individual medley and set a world record in a time of 4 minutes 8.26 seconds. Panacea [classification: correct]: Sports
Appendix E Implementation of Baselines